# IterResearch: Rethinking Long-Horizon Agents via Markovian State Reconstruction

### v1

@@ Appendix D Case Study of IterResearch
I need to understand the biology of Micromalthidae. I recall that Micromalthidae is a family of beetles, and they might have unique feeding habits or life cycles. Specifically, I think Micromalthus debilis is a species in this family, and it’s known for having a complex life cycle with paedogenesis, where larvae can produce offspring. But the question is about an adult male and what it fed on. Adult males might not feed at all if they have a short lifespan or if their role is solely reproduction. The answer choices include "Its mother," which sounds bizarre, but it might be related to matriphagy or something similar. Also, fungus and decaying wood are common food sources for wood-feeding insects, but cellulose is a component of wood. I should verify the feeding habits of Micromalthidae adults. Since this is a specific entomological question, I need to search for reliable information. Available tools include google_search, google_scholar, Visit, and PythonInterpreter. For this, I should use google_scholar or google_search to find academic sources on Micromalthidae feeding behavior. Let me plan my approach:
@@
I remember that in some insects, adult males don’t feed, but I need confirmation for this family. Also, the question mentions "upon its death," which might imply that it didn’t feed much or at all, leading to choice E. Nothing. But let’s not jump to conclusions. I need evidence. I’$l$ start by searching using google_scholar for Micromalthidae feeding habits or life cycle. Since the question is in English, my output should be in English. Now, for the report section, I need to provide: - Problem Essence Analysis: Deeply analyze core elements, complexity levels, and potential pitfalls. - Problem-Solving Planning: Based on understanding, clarify what needs to be confirmed, information sufficiency, and cognitive blind spots. Then, I must make one tool call. Potential pitfalls: Misinterpreting the life cycle or feeding habits. Micromalthidae might have unusual behaviors. Cognitive blind spots: Assuming adult beetles always feed, but some might not. Information sufficiency: I currently don’t know enough, so I need to search. I should search for "Micromalthidae adult feeding" or "Micromalthus debilis life cycle". Let me formulate the search query. Query: "Micromalthidae adult male feeding" Or more broadly: "Micromalthidae life cycle" But the question specifies adult male, so perhaps focus on that. Also, "matriphagy in Micromalthidae" because choice A is "Its mother," which might relate to that. I recall that in Micromalthus debilis, there is matriphagy, where larvae eat the mother. But the question is about an adult male, not larva. The adult male might not feed, or perhaps it has a different diet. Upon second thought, in some sources, adult Micromalthus might not feed at all. But let’s verify. I’$l$ use google_scholar to search. For the tool call, I need to specify the query. Let me decide on the query. Since google_scholar returns academic papers, I should use precise terms. Query: "Micromalthidae adult feeding habits" Or "Micromalthus debilis adult nutrition" But to be safe, I’$l$ use multiple queries to cover bases. The tool allows multiple queries, so I can do an array. Let me think. First, I need to confirm the feeding behavior. Also, the Visit tool might be useful if I find specific URLs, but first, I need to search. PythonInterpreter probably isn’t needed for this, as it’s about information retrieval, not calculation. So, I’$l$ use google_scholar. Now, for the report section. Problem Essence Analysis: The core is to determine the diet of an adult male Micromalthidae insect based on its life cycle and biological characteristics. Complexity arises from potential unique behaviors like matriphagy or non-feeding adults. Pitfalls include misattributing larval behaviors to adults or generalizing from other beetles. Problem-Solving Planning: I need to confirm the feeding habits of adult males in Micromalthidae. Current knowledge is insufficient, so I must search academic sources. Blind spots might be assuming all adults feed or not, without specific data. Then, tool call: Use google_scholar with queries about Micromalthidae adult feeding. Let me write the queries. Queries: ["Micromalthidae adult male feeding", "Micromalthus debilis life cycle", "matriphagy Micromalthidae"] I included matriphagy because it might be relevant, even though it’s about larvae eating mother, but perhaps adults have connections. But the question is about adult male feeding, so focus on that. Perhaps "do Micromalthidae adults feed?" But let’s go with the queries. Now, output format: First <<report>>, then <<tool_call>>. Language: English. So, draft the report. <</think>>
@@
"name": "google_scholar", "arguments": "query": ["Micromalthidae adult male feeding", "Micromalthus debilis life cycle", "matriphagy Micromalthidae", "do Micromalthidae adults feed"]
@@ A Google scholar for ’Micromalthidae adult male feeding’ found 10 results:
Looking at the previous tool call, I used Google Scholar to search for information on Micromalthidae feeding habits, life cycle, and matriphagy. The results from the tool response are quite detailed. Several sources mention that male larvae practice matriphagy, eating their mothers. For example, the Current Biology article by Normark (2013) states that it takes a male about a week to devour his entire mother before pupating into an adult male. Other papers like Pollock & Normark (2002) and Perotti et al. (2016) also confirm this behavior. The Wikipedia entry and other web results reinforce this, mentioning that adult males rarely develop and consume their mothers. The key point here is that adult males themselves don’t feed; it’s the male larvae that engage in matriphagy. Once they become adults, they might not feed at all, as indicated by references to vestigial reproductive organs and non-functional adults. So, putting it together: the adult male hasn’t fed anything after becoming an adult because its feeding happened during the larval stage as part of matriphagy. Therefore, the only thing it fed on was its mother. The correct answer should be A. Its mother. I need to make sure there’s no conflicting info. Some sources mention larvae feeding on decaying wood or fungus, but specifically for males, the matriphagy is highlighted. Also, the question specifies "upon its death," referring to the adult male, so we’re looking back at what it consumed during its life. Since the adult doesn’t feed, the last thing it ate was its mother during the larval stage. The status report should document all findings, emphasizing the matriphagy behavior and the fact that adult males don’t feed. The next steps aren’t needed since the information is sufficient from the sources. So, the final answer is A. <</think>>
@@
**Progress Achieved:** Based on comprehensive research through Google Scholar and web searches, I have gathered substantial information about Micromalthidae biology, specifically focusing on Micromalthus debilis, the only extant species in this family. The key findings are: 1. **Unique Life Cycle Characteristics**: Micromalthus debilis exhibits one of the most complex life cycles among beetles, involving multiple reproductive strategies including paedogenesis (reproduction by larvae), haplodiploidy, thelytoky, and arrhenotoky. The species is primarily parthenogenetic with female larvae reproducing asexually. 2. **Male Development and Feeding Behavior**: Male development occurs through a specialized pathway where male eggs hatch into active triungulin larvae that immediately begin feeding. Multiple authoritative sources (Normark 2013 in Current Biology, Pollock & Normark 2002, Perotti et al. 2016) confirm that male larvae exhibit obligate matriphagy - they consume their mothers from within. According to Normark (2013): "It takes a male about a week to devour his entire mother, after which he pupates and molts into a winged adult male."
@@
3. **Adult Feeding Habits**: Adult males of Micromalthus debilis appear to be non-feeding or have vestigial mouthparts. Research by Beutel & Hörnschemeyer (2002) and Yavorskaya et al. (2018) indicates that adults have reduced morphological features related to feeding, suggesting they do not feed as adults. The Wikipedia entry on telephone-pole beetles confirms that adult beetles have "vestigial reproductive organs" and likely do not feed. 4. **Contrast with Larval Feeding**: While larvae (particularly female larvae) do feed on decaying wood and associated fungi, this feeding behavior is specific to the larval stage and different between sexes. Male larvae have a specialized feeding requirement focused on matriphagy rather than wood consumption. 5. **Ecological Context**: The laboratory colony described in the question provides decaying wood, but this substrate is primarily utilized by female larvae rather than male individuals. The matriphagy behavior represents an evolutionary adaptation to nutrient acquisition in nutrient-poor environments. **Source Verification Status**: - High confidence in Normark (2013) Current Biology article and Pollock & Normark (2002) journal paper - Supporting evidence from Perotti et al. (2016) Scientific Reports - Morphological studies (Beutel & Hörnschemeyer 2002, Yavorskaya et al. 2018) provide additional confirmation - Wikipedia and other secondary sources align with academic literature **Uncertainties Resolved**: - Confirmed that adult males do not feed as adults - Established that male larvae exclusively consume their mothers - Clarified that decaying wood feeding is primarily associated with female larvae **Next Steps Plan:** No further research is required as sufficient authoritative information has been obtained to answer the question definitively. The collected evidence clearly establishes the feeding history of adult male Micromalthidae insects. <</report>>
@@ E.2 Instruction of our IterResearch
Prompt of our IterResearch You are a professional problem-solving agent with rigorous information verification capabilities and deep analytical thinking. ## CRITICAL OUTPUT FORMAT REQUIREMENTS You MUST follow this exact format. Every response must contain: 1. <<report>><</report>> (always required) 2. Either <<answer>><</answer>> OR <<tool_call>><</tool_call>> (never both) ## Input Format - **Current Date**: Current Date - **Question**: The problem posed by the user that needs to be solved - **Last Status Report and Deep Analysis**: A summary overview of current work progress - **Last Tool Call**: The specific action taken in the previous round - **Last Observation**: The results and feedback obtained after the previous action ## Output Format <<report>> ### Status Report and Deep Analysis **Progress Achieved:** Based on the Last Status Report and Deep Analysis and Last Tool Response provided in the input, compile a comprehensive and complete documentation of all currently collected information, conclusions, data, and findings. This section must capture ALL important information without any omissions, presented in plain text format with corresponding sources clearly annotated. You must directly record the actual information content rather than using referential markers or summaries. This includes: 1. All factual data and evidence collected 2. All analytical conclusions and insights derived 3. All source materials and their verification status 4. All uncertainties, limitations, or gaps identified 5. Complete integration of previous progress with new findings The documentation must be sufficiently detailed and complete that someone can fully inherit and understand all achieved progress to seamlessly continue the research without losing any critical information or context. **Next Steps Plan:** Based on the comprehensive progress achieved above, formulate a detailed and actionable plan for the next phase of research or investigation. <</report>> You MUST output this section enclosed with <<report>><</report>> tags! **Decision Point**: Are you certain that no further verification or information gathering is needed to provide the final answer? **If YES - Information is sufficient:** <<answer>> Answer Format: 1. **Language**: Your answer should be in the same language as the question. If the question uses English, answer in English. If the question uses Chinese, answer in Chinese. 2. The answer should include as much relevant content as possible. Organize the content into separate paragraphs to avoid overly long sections. Avoid content duplication in the answer. 3. Do not include any non-text elements such as URLs, images, or tables that appeared in the reasoning. 4. Output only the answer text. Do not use any additional symbols or start with phrases like ’Here is my answer’. 5. First, output a direct answer to the question. 6. Do not just output the answer to the question; provide a rich and lengthy response by synthesizing all relevant information, and format it using markdown. 7. For statistical data with at least 3 items, use a markdown table to present the results, ensuring the table description is clear. For less than 3 items, describe them directly in text. 8. For research-type questions, try to generate a report of over 1000 words, using subheadings and other elements to improve readability and logic. <</answer>> You MUST output this section enclosed with <<answer>><</answer>> tags! **If NO - Further action needed:** <<tool_call>> "name": "tool name here", "arguments": "parameter name here": parameter value here, "another parameter name here": another parameter value here, … <</tool_call>> You MUST output this section enclosed with <<tool_call>><</tool_call>> tags! ## Working Principles 1. **Rigorous Verification**: Critically evaluate all information sources 2. **Deep Thinking**: Pursue essential understanding, not satisfied with surface phenomena 3. **Evidence-Driven**: Make reasoning decisions based on reliable evidence through deep thinking 4. **You are required to maintain detailed documentation in all your reports and actions, providing sufficient information for others to fully grasp your progress and effectively continue or modify the research trajectory based on your contributions.** ## Special Requirements - All tools in the tool list are real and functional - as long as you make correct tool calls, you will receive their returned results. - Clearly distinguish between "confirmed facts," "highly credible inferences," and "hypotheses to be verified" - Clearly indicate uncertainty when information is insufficient - Always focus on the original question - When outputting [Status Report and Deep Analysis], never omit key actions and results, even if these actions or results do not meet expectations, these conclusions must still be documented. - **When further action is needed, you must select an appropriate tool from your available tool list and carefully configure the tool call parameters based on the tool’s specific characteristics and requirements** - **When the current status is sufficient to answer the question, must provide the final answer enclosed with <<answer>><</answer>> tags rather than continue with actions** ## FORMAT REMINDER - Start with <<report>><</report>> section - Then choose: <<answer>><</answer>> if sufficient info, OR <<tool_call>><</tool_call>> if need more action - Never output both answer and tool_call tags in same response ## Input - Current Date: {date_to_use}

### v2

@@ Appendix D Case Study of IterResearch
I need to understand the biology of Micromalthidae. I recall that Micromalthidae is a family of beetles, and they might have unique feeding habits or life cycles. Specifically, I think Micromalthus debilis is a species in this family, and it’s known for having a complex life cycle with paedogenesis, where larvae can produce offspring. But the question is about an adult male and what it fed on. Adult males might not feed at all if they have a short lifespan or if their role is solely reproduction. The answer choices include ”Its mother,” which sounds bizarre, but it might be related to matriphagy or something similar. Also, fungus and decaying wood are common food sources for wood-feeding insects, but cellulose is a component of wood. I should verify the feeding habits of Micromalthidae adults. Since this is a specific entomological question, I need to search for reliable information. Available tools include google_search, google_scholar, Visit, and PythonInterpreter. For this, I should use google_scholar or google_search to find academic sources on Micromalthidae feeding behavior. Let me plan my approach:
@@
I remember that in some insects, adult males don’t feed, but I need confirmation for this family. Also, the question mentions ”upon its death,” which might imply that it didn’t feed much or at all, leading to choice E. Nothing. But let’s not jump to conclusions. I need evidence. I’$l$ start by searching using google_scholar for Micromalthidae feeding habits or life cycle. Since the question is in English, my output should be in English. Now, for the report section, I need to provide: - Problem Essence Analysis: Deeply analyze core elements, complexity levels, and potential pitfalls. - Problem-Solving Planning: Based on understanding, clarify what needs to be confirmed, information sufficiency, and cognitive blind spots. Then, I must make one tool call. Potential pitfalls: Misinterpreting the life cycle or feeding habits. Micromalthidae might have unusual behaviors. Cognitive blind spots: Assuming adult beetles always feed, but some might not. Information sufficiency: I currently don’t know enough, so I need to search. I should search for ”Micromalthidae adult feeding” or ”Micromalthus debilis life cycle”. Let me formulate the search query. Query: ”Micromalthidae adult male feeding” Or more broadly: ”Micromalthidae life cycle” But the question specifies adult male, so perhaps focus on that. Also, ”matriphagy in Micromalthidae” because choice A is ”Its mother,” which might relate to that. I recall that in Micromalthus debilis, there is matriphagy, where larvae eat the mother. But the question is about an adult male, not larva. The adult male might not feed, or perhaps it has a different diet. Upon second thought, in some sources, adult Micromalthus might not feed at all. But let’s verify. I’$l$ use google_scholar to search. For the tool call, I need to specify the query. Let me decide on the query. Since google_scholar returns academic papers, I should use precise terms. Query: ”Micromalthidae adult feeding habits” Or ”Micromalthus debilis adult nutrition” But to be safe, I’$l$ use multiple queries to cover bases. The tool allows multiple queries, so I can do an array. Let me think. First, I need to confirm the feeding behavior. Also, the Visit tool might be useful if I find specific URLs, but first, I need to search. PythonInterpreter probably isn’t needed for this, as it’s about information retrieval, not calculation. So, I’$l$ use google_scholar. Now, for the report section. Problem Essence Analysis: The core is to determine the diet of an adult male Micromalthidae insect based on its life cycle and biological characteristics. Complexity arises from potential unique behaviors like matriphagy or non-feeding adults. Pitfalls include misattributing larval behaviors to adults or generalizing from other beetles. Problem-Solving Planning: I need to confirm the feeding habits of adult males in Micromalthidae. Current knowledge is insufficient, so I must search academic sources. Blind spots might be assuming all adults feed or not, without specific data. Then, tool call: Use google_scholar with queries about Micromalthidae adult feeding. Let me write the queries. Queries: [”Micromalthidae adult male feeding”, ”Micromalthus debilis life cycle”, ”matriphagy Micromalthidae”] I included matriphagy because it might be relevant, even though it’s about larvae eating mother, but perhaps adults have connections. But the question is about adult male feeding, so focus on that. Perhaps ”do Micromalthidae adults feed?” But let’s go with the queries. Now, output format: First <<report>>, then <<tool_call>>. Language: English. So, draft the report. <</think>>
@@
”name”: ”google_scholar”, ”arguments”: ”query”: [”Micromalthidae adult male feeding”, ”Micromalthus debilis life cycle”, ”matriphagy Micromalthidae”, ”do Micromalthidae adults feed”]
@@ A Google scholar for ’Micromalthidae adult male feeding’ found 10 results:
Looking at the previous tool call, I used Google Scholar to search for information on Micromalthidae feeding habits, life cycle, and matriphagy. The results from the tool response are quite detailed. Several sources mention that male larvae practice matriphagy, eating their mothers. For example, the Current Biology article by Normark (2013) states that it takes a male about a week to devour his entire mother before pupating into an adult male. Other papers like Pollock & Normark (2002) and Perotti et al. (2016) also confirm this behavior. The Wikipedia entry and other web results reinforce this, mentioning that adult males rarely develop and consume their mothers. The key point here is that adult males themselves don’t feed; it’s the male larvae that engage in matriphagy. Once they become adults, they might not feed at all, as indicated by references to vestigial reproductive organs and non-functional adults. So, putting it together: the adult male hasn’t fed anything after becoming an adult because its feeding happened during the larval stage as part of matriphagy. Therefore, the only thing it fed on was its mother. The correct answer should be A. Its mother. I need to make sure there’s no conflicting info. Some sources mention larvae feeding on decaying wood or fungus, but specifically for males, the matriphagy is highlighted. Also, the question specifies ”upon its death,” referring to the adult male, so we’re looking back at what it consumed during its life. Since the adult doesn’t feed, the last thing it ate was its mother during the larval stage. The status report should document all findings, emphasizing the matriphagy behavior and the fact that adult males don’t feed. The next steps aren’t needed since the information is sufficient from the sources. So, the final answer is A. <</think>>
@@
**Progress Achieved:** Based on comprehensive research through Google Scholar and web searches, I have gathered substantial information about Micromalthidae biology, specifically focusing on Micromalthus debilis, the only extant species in this family. The key findings are: 1. **Unique Life Cycle Characteristics**: Micromalthus debilis exhibits one of the most complex life cycles among beetles, involving multiple reproductive strategies including paedogenesis (reproduction by larvae), haplodiploidy, thelytoky, and arrhenotoky. The species is primarily parthenogenetic with female larvae reproducing asexually. 2. **Male Development and Feeding Behavior**: Male development occurs through a specialized pathway where male eggs hatch into active triungulin larvae that immediately begin feeding. Multiple authoritative sources (Normark 2013 in Current Biology, Pollock & Normark 2002, Perotti et al. 2016) confirm that male larvae exhibit obligate matriphagy - they consume their mothers from within. According to Normark (2013): ”It takes a male about a week to devour his entire mother, after which he pupates and molts into a winged adult male.”
@@
3. **Adult Feeding Habits**: Adult males of Micromalthus debilis appear to be non-feeding or have vestigial mouthparts. Research by Beutel & Hörnschemeyer (2002) and Yavorskaya et al. (2018) indicates that adults have reduced morphological features related to feeding, suggesting they do not feed as adults. The Wikipedia entry on telephone-pole beetles confirms that adult beetles have ”vestigial reproductive organs” and likely do not feed. 4. **Contrast with Larval Feeding**: While larvae (particularly female larvae) do feed on decaying wood and associated fungi, this feeding behavior is specific to the larval stage and different between sexes. Male larvae have a specialized feeding requirement focused on matriphagy rather than wood consumption. 5. **Ecological Context**: The laboratory colony described in the question provides decaying wood, but this substrate is primarily utilized by female larvae rather than male individuals. The matriphagy behavior represents an evolutionary adaptation to nutrient acquisition in nutrient-poor environments. **Source Verification Status**: - High confidence in Normark (2013) Current Biology article and Pollock & Normark (2002) journal paper - Supporting evidence from Perotti et al. (2016) Scientific Reports - Morphological studies (Beutel & Hörnschemeyer 2002, Yavorskaya et al. 2018) provide additional confirmation - Wikipedia and other secondary sources align with academic literature **Uncertainties Resolved**: - Confirmed that adult males do not feed as adults - Established that male larvae exclusively consume their mothers - Clarified that decaying wood feeding is primarily associated with female larvae **Next Steps Plan:** No further research is required as sufficient authoritative information has been obtained to answer the question definitively. The collected evidence clearly establishes the feeding history of adult male Micromalthidae insects. <</report>>
@@ E.2 Instruction of our IterResearch
Prompt of our IterResearch You are a professional problem-solving agent with rigorous information verification capabilities and deep analytical thinking. ## CRITICAL OUTPUT FORMAT REQUIREMENTS You MUST follow this exact format. Every response must contain: 1. <<report>><</report>> (always required) 2. Either <<answer>><</answer>> OR <<tool_call>><</tool_call>> (never both) ## Input Format - **Current Date**: Current Date - **Question**: The problem posed by the user that needs to be solved - **Last Status Report and Deep Analysis**: A summary overview of current work progress - **Last Tool Call**: The specific action taken in the previous round - **Last Observation**: The results and feedback obtained after the previous action ## Output Format <<report>> ### Status Report and Deep Analysis **Progress Achieved:** Based on the Last Status Report and Deep Analysis and Last Tool Response provided in the input, compile a comprehensive and complete documentation of all currently collected information, conclusions, data, and findings. This section must capture ALL important information without any omissions, presented in plain text format with corresponding sources clearly annotated. You must directly record the actual information content rather than using referential markers or summaries. This includes: 1. All factual data and evidence collected 2. All analytical conclusions and insights derived 3. All source materials and their verification status 4. All uncertainties, limitations, or gaps identified 5. Complete integration of previous progress with new findings The documentation must be sufficiently detailed and complete that someone can fully inherit and understand all achieved progress to seamlessly continue the research without losing any critical information or context. **Next Steps Plan:** Based on the comprehensive progress achieved above, formulate a detailed and actionable plan for the next phase of research or investigation. <</report>> You MUST output this section enclosed with <<report>><</report>> tags! **Decision Point**: Are you certain that no further verification or information gathering is needed to provide the final answer? **If YES - Information is sufficient:** <<answer>> Answer Format: 1. **Language**: Your answer should be in the same language as the question. If the question uses English, answer in English. If the question uses Chinese, answer in Chinese. 2. The answer should include as much relevant content as possible. Organize the content into separate paragraphs to avoid overly long sections. Avoid content duplication in the answer. 3. Do not include any non-text elements such as URLs, images, or tables that appeared in the reasoning. 4. Output only the answer text. Do not use any additional symbols or start with phrases like ’Here is my answer’. 5. First, output a direct answer to the question. 6. Do not just output the answer to the question; provide a rich and lengthy response by synthesizing all relevant information, and format it using markdown. 7. For statistical data with at least 3 items, use a markdown table to present the results, ensuring the table description is clear. For less than 3 items, describe them directly in text. 8. For research-type questions, try to generate a report of over 1000 words, using subheadings and other elements to improve readability and logic. <</answer>> You MUST output this section enclosed with <<answer>><</answer>> tags! **If NO - Further action needed:** <<tool_call>> ”name”: ”tool name here”, ”arguments”: ”parameter name here”: parameter value here, ”another parameter name here”: another parameter value here, … <</tool_call>> You MUST output this section enclosed with <<tool_call>><</tool_call>> tags! ## Working Principles 1. **Rigorous Verification**: Critically evaluate all information sources 2. **Deep Thinking**: Pursue essential understanding, not satisfied with surface phenomena 3. **Evidence-Driven**: Make reasoning decisions based on reliable evidence through deep thinking 4. **You are required to maintain detailed documentation in all your reports and actions, providing sufficient information for others to fully grasp your progress and effectively continue or modify the research trajectory based on your contributions.** ## Special Requirements - All tools in the tool list are real and functional - as long as you make correct tool calls, you will receive their returned results. - Clearly distinguish between ”confirmed facts,” ”highly credible inferences,” and ”hypotheses to be verified” - Clearly indicate uncertainty when information is insufficient - Always focus on the original question - When outputting [Status Report and Deep Analysis], never omit key actions and results, even if these actions or results do not meet expectations, these conclusions must still be documented. - **When further action is needed, you must select an appropriate tool from your available tool list and carefully configure the tool call parameters based on the tool’s specific characteristics and requirements** - **When the current status is sufficient to answer the question, must provide the final answer enclosed with <<answer>><</answer>> tags rather than continue with actions** ## FORMAT REMINDER - Start with <<report>><</report>> section - Then choose: <<answer>><</answer>> if sufficient info, OR <<tool_call>><</tool_call>> if need more action - Never output both answer and tool_call tags in same response ## Input - Current Date: {date_to_use}